OpenStudy (anonymous):

Which value of x satisfies both 3x + y = 9 and 3x − 5y = 15?

OpenStudy (anonymous):

\[\frac{ 10 }{ 3 }\] \[\frac{ 2 }{ 3}\] -1 0

OpenStudy (mrnood):

subtract the 2 equations - this will eliminate x and leave you with an equation for y which you can solve. Then use THAT value of y in one of eh equations to get x

OpenStudy (anonymous):

Well, let's set this up using elimination. 3x+y=9 Multiply everything by 5 5(3x+y=9) 15x+5y=45 + 3x-5y=15 ---This is the second equation, we are eliminatating y. 18x=60 x = 60/18 x=30/9 x=10/9 And there you go!

OpenStudy (anonymous):

Whoops, the last one is supposed to be x = 10/3, sorry about that.

OpenStudy (mrnood):

OR Just blag the answer that someone else gives you without having to think for yourself at all You choose.

OpenStudy (anonymous):

What??@MrNood

OpenStudy (mrnood):

It is against the code of conduct ( @NinjaDevo ) to post complete answers. YOU can take his answer and post it on you multichoice site or paper - but you will not learn anything by doing so, as the answer to THIS question is not the answer to the next. Do you understand the concept of using elimination to solve this type of ' equation system' or 'simultaneous equations?

OpenStudy (anonymous):

Oh...Sorry I'm a little new around here. I didn't know that was against the code of conduct. :o I was trying to show him how to solve the equation so hopefully he could solve one similar to it in the future. Sorry for all the confusion guys, this is probably my fault.

OpenStudy (mrnood):

Ok - soz if I sound a little hard on you. Thx for contributing. I believe it is better if you leave the original questioner to solve the question, based on concepts and hints you may give. This gives them a method to apply to OTHER questions, rather than just a quick fix for the current one.

OpenStudy (anonymous):

Alright, thanks MrNood. I'll try to do that in the future.